OpenStudy (anonymous):

Solve x^2 - 9 = 0 by factoring.

OpenStudy (anonymous):

Using difference of squares, you get (x-3)(x+3), Therefore the roots are 3 and negative three.

OpenStudy (anonymous):

(x-3)(x+3) x=3 or x=-3

OpenStudy (anonymous):

What about this one? i'm lost on it Solve x2 + 5x + 6 = 0 using factoring.

OpenStudy (anonymous):

(x+3)(x+2)=0 x=-3 or x=-2

OpenStudy (anonymous):

Where did you get the 3

OpenStudy (anonymous):

Because 2*3=6 and 2+3=5, so it factors into (x+3)(x+2). And when the equation is factored out, the only two numbers that are true are the ones that make one factor or the other equal zero so the whole thing does and the statement is true.

OpenStudy (anonymous):

Thank you